# A Survey on All-in-One Image Restoration: Taxonomy, Evaluation and Future Trends

### v2

@@ Abstract
Image restoration (IR) aims to recover high-quality images from inputs degraded by various factors such as noise, blur, compression, and adverse weather. Traditional IR methods typically focus on specific types of degradation, which limits their effectiveness in real-world scenarios with complex distortions. In response to this challenge, the all-in-one image restoration (AiOIR) paradigm has recently emerged, offering a unified framework that adeptly addresses multiple degradation types. These innovative models enhance convenience and versatility by adaptively learning degradation-specific features while simultaneously leveraging shared knowledge across diverse corruptions. In this survey, we present the first comprehensive overview of AiOIR, offering a taxonomy that organizes existing methods by architecture innovations, learning strategies, and key improvements. We systematically categorize prevailing approaches and critically assess the challenges these models encounter, proposing future research directions to propel this rapidly evolving field. Our survey begins with an introduction to the foundational concepts of AiOIR models, followed by a categorization of typical scenarios. We then highlight key architectural and algorithmic advances in AiOIR, aiming to inspire continued innovation. To facilitate rigorous evaluation of existing methods, we collate and summarize established datasets, evaluation metrics, and common experimental settings. Finally, we present an objective comparison of open-sourced methods, providing valuable insights for researchers and practitioners. This paper stands as the first comprehensive and insightful review of all-in-one image restoration. A related repository is available at https://github.com/Harbinzzy/All-in-One-Image-Restoration-Survey.
@@ I Introduction
Image processing is an integral part of low-level vision tasks, and digital image processing has evolved significantly over the past few decades, transitioning from traditional methods to advanced deep learning techniques. These methods, while effective, are limited by their inability to handle complex and varied image degradation scenarios. With the rise of deep learning, image processing has achieved remarkable results, especially driven by convolutional neural networks (CNNs) [1], Transformers [2]. In the image restoration (IR) field, single-task IR has achieved notable breakthroughs in correcting a specific type of image degradation (e.g., denoising [3, 4, 5, 6, 7, 8], dehazing [9, 10, 11, 12, 13], desnowing [14, 15, 16], deraining [17, 18, 19], deblurring [20, 21] and low-light image enhancement [22, 23, 24]). Although they have set state-of-the-art performance on restoring diverse degraded images, existing single-task IR methods generally lack the flexibility to adapt to new types of image degradation without extensive retraining (it is worth noting that some real-world IR methods [25, 26, 27, 28] are not categorized as single-task IR). To alleviate the aforementioned issues, researchers have proposed all-in-one image restoration (AiOIR), devoting significant effort to introduce various key improvements (e.g., prompt-based learning approach, mixture-of-experts structure, multimodal model and autonomous agent) beyond single-task IR.
@@
With the rapid development of AiOIR, numerous researchers have collected a series of datasets tailored for various all-in-one task settings, e.g., the All-Weather dataset [29] for synthetic weather-related scenarios, the WeatherStream dataset [30] for real-world weather-related scenarios, the CDD-11 dataset [31] for composite degradations, etc. Leveraging these datasets, most recent works focus on improving the representation capability of IR networks for complicated degradation through well-designed methods based on prompt learning, contrastive learning, multimodal representations, etc. Although these works have achieved superior progress in the objective quality(e.g., PSNR, NIQE [32] and FID [33]), the restored images still suffer from unsatisfactory texture generation, hindering the application of IR methods in real-world scenarios. Collectively, these methodologies represent a substantial progression in the pursuit of sophisticated, accurate, and versatile AiOIR solutions.
@@
    Methodology (Sec. III) : We provide a detailed analysis of representative AiOIR networks, aiming to illustrate the prevailing methods and elucidate the different categories of methods. By analyzing state-of-the-art methods, we summarize network architectures, learning strategies, and some key improvements.
@@ II Preliminaries
In this section, we first define the concept of all-in-one image restoration and compare it with related concepts. Subsequently, we provide a comprehensive review of diverse scenarios in AiOIR and the corresponding approaches. Finally, we illustrate the interactions between different IR tasks.
@@ II-A Concept Definitions
Single degradation image restoration (SDIR) focuses on recovering clean images from observations that have been corrupted by a specific type of degradation, such as noise [4, 5, 6], blur [20, 21], or rain [17, 18, 19]. These methods are typically task-specific, meaning that each model is designed for a particular degradation type. While these models demonstrate strong performance on known degradation types, they often exhibit limited generalization capability when encountering unseen degradation patterns or novel corruption intensities. Recently, some works [50, 51, 52] show certain generalizability to various degradation types with a unified framework. However, they need to train different models for different degradations, which are not all-in-one solutions as expected in practice. We refer to these methods as multiple degradation image restoration (MDIR). Subsequently, there have been recent attempts to unify multiple image restoration tasks within a single framework without the need for retraining. The seminal work [53] proposed the all-in-one weather restoration method which utilizes a multi-encoder and decoder architecture and the neural architecture search across task-specific optimized encoders. Moreover, Li et al. [54] first proposed the all-in-one blind image restoration using contrastive learning to extract degradation representations, which can recover images from multiple corruptions without any degradation information in advance. The AiOIR approach aims to recover clean images under various conditions explicitly tailored to address isolated or mixed degradation based on multi-head and multi-tail structures, priors, or pretrained models within a unified framework. It is worth noting that, some robust IR models [25, 27, 26, 28, 55, 56] are still broadly considered part of the AiOIR setting. The AiOIR models present practical advantages such as reduced storage requirements and simplified deployment. However, the principal challenge lies in developing robust architectures capable of effectively addressing diverse degradations and achieving high-quality restorations across varying conditions in real-world scenarios.
Figure 2: Milestones of AiOIR methods. We simply list their names and venues. The bar chart shows the approximate annual number of papers on AiOIR from 2020 to May 2025.
@@ II-B Typical Scenarios
We categorize AiOIR into several scenarios according to types of images processed, including natural images, images under adverse weather conditions, underwater images, medical images, document images, and hyperspectral images.
Natural images. The degradation types commonly encountered in natural images include Gaussian noise, real-world noise, defocus blur, motion blur, low-light conditions, JPEG compression artifacts, mosaic effects, and issues arising from under-display cameras. Various AiOIR approaches have been developed to address these degradations, including techniques outlined in studies such as [57, 58, 59, 60, 61, 62]. These approaches primarily focus on resolving typical problems found in both natural and man-made scenes, enhancing the quality and usability of images affected by these common degradations.
Adverse weather conditions. In more extreme cases, the images requiring restoration may be severely impacted by various adverse weather conditions, such as snowflakes, raindrops, and dense haze effects. These conditions lead to ill-posed inverse problems that are crucial for applications in autonomous navigation, video retrieval, and outdoor surveillance. The AiOIR field has seen the emergence of many solutions aiming at addressing these challenges, including works like [63, 29, 64, 65, 66, 67]. Notably, Liu et al. [68] proposed multi-weather nighttime image restoration. These solutions are designed to effectively restore visibility and clarity in images affected by harsh environmental conditions.
Underwater images. Underwater images suffer from wavelength-dependent light attenuation and scattering by marine microorganisms, leading to haze, color casts, low contrast, and loss of fine detail. UniUIR [56], a pioneering framework to tackle such compound distortions in an integrated manner, introduces the Mamba MoE module leveraging DepthAnythingV2 [69] to inject scene depth priors.
@@
Document images. DocRes [72] pioneers the exploration of generalist models for all-in-one restoration of document images. This work addresses five tasks (i.e., dewarping, deshadowing, appearance enhancement, deblurring, and binarization). DocRes employs a straightforward yet highly effective visual prompting method known as DTSPrompt, which effectively distinguishes between different tasks and accommodates various resolutions. TextDoctor [73], utilizing the structure pyramid prediction and patch pyramid diffusion model, exhibits superior performance in capturing the structure of inpainted text.
@@ II-C Task Interactions
These intertwined degradations present both challenges (task entanglement, ordering sensitivity, domain ambiguity) and opportunities (shared features, multi-task learning, unified representations) for AiOIR models. Understanding these interdependencies is crucial for the design of robust and generalizable AiOIR systems, especially in unconstrained or real-world environments where degradation types and combinations are unpredictable.
@@ III Methodology
All-in-one image restoration methods have gained significant attention for their ability to address multiple types of degradation within a unified model. These methods are particularly valuable in real-world scenarios where images may be simultaneously affected by various artifacts, such as noise, blur, and adverse weather conditions. This section presents a comprehensive analysis of AiOIR methods from multiple perspectives. Firstly, we examine both novel and traditional network architectures, discussing their implementation approaches and classifying them into several typical network designs. Subsequently, we further investigate diverse learning strategies, which encompass training methodologies and perspectives, to improve image reconstruction accuracy and overall restoration quality. Finally, we examine other advanced techniques that optimize AiOIR models, including Prompt Learning, Mixture-of-Experts (MoE), Multimodal Model, and Agentic IR System. In Fig. 2, we present the representative works on AiOIR according to the timeline and method types.
@@ III-A Network Design for AiOIR
Existing all-in-one image restoration methods, as a form of multi-task learning (MTL), employ various architectural designs to handle inputs and outputs across multiple tasks, enabling efficient information sharing among them. Despite significant differences in their architectures, AiOIR methods can be broadly categorized into ten representative frameworks, as illustrated in Fig. 3 and described below.
@@
(f) Mixture-of-Experts Architecture: In MoE-based models, inputs are routed to different experts via a gating mechanism. For instance, MEASNet [82] considers both pixel-level and global features (including low-frequency and high-frequency components) to select the appropriate expert for image restoration.
@@
(h) Visual Prompting: These models employ a single encoder-decoder architecture and inject learnable visual prompts at multiple decoding stages to implicitly predict degradation conditions. Examples include works like [58, 89, 90, 91, 92]. The prompts guide the decoder in adaptively recovering various degraded images, serving as lightweight plug-and-play modules with minimal additional parameters.
@@ III-B Learning Strategies of AiOIR
Besides network design (as discussed in Sec. III-A), robust learning strategies are crucial for achieving satisfactory results in AiOIR. In this section, we discuss several promising learning strategies in this field. We begin with continual learning, exploring how it can prevent catastrophic forgetting. Next, we focus on contrastive learning and its application in complex degradation scenarios, emphasizing its role in enhancing the discriminative power of image features. We then discuss the zero-shot learning in AiOIR, emphasizing its ability to generalize to unseen degradations. Additionally, we highlight the potential of multi-task learning (MTL) to optimize performance across various degradation tasks, pointing out the importance of addressing task relationships and conflicts. Finally, we introduce the concept of machine unlearning, exploring its potential for privacy protection.
@@ III-B2 Contrastive Learning
One of the significant challenges in IR is effectively handling unseen tasks and degradation types. The vast variability in possible degradations can severely impede a model’s generalization capabilities, making it less effective when confronted with new, unseen data. To address this issue, researchers have drawn inspiration from contrastive learning techniques that have proven successful in both high-level and low-level tasks [109, 110, 111]. Contrastive learning approaches typically serve as an additional form of regularization to improve the generalization of single-task restoration models [112, 113, 111, 114, 115]. By incorporating contrastive regularization, these methods aim to enhance model performance across a wide array of image restoration applications. In contrastive learning, the definitions of positive and negative samples can be flexibly adjusted, allowing researchers to tailor the learning process to better suit specific tasks and datasets. This flexibility ultimately enhances the model’s adaptability and performance in diverse image restoration scenarios [84]. Moreover, contrastive learning-based loss functions have been proposed to obtain discriminative degradation representations in the latent space [54, 57, 89, 116], further improving the AiOIR model’s ability to distinguish between different types of degradations and to generalize to unseen ones.
@@ III-B4 Multi-task Learning
MTL is a learning paradigm that leverages shared representations and knowledge across tasks, allowing models to learn more efficiently and effectively. By jointly learning from multiple objectives, MTL can improve generalization, reduce overfitting, and achieve better performance on individual tasks. Traditional MTL typically involves a combination of semantic tasks (e.g., classification, detection, segmentation). The tasks exhibit significant semantic differences and also possess distinct output spaces. Typical MTL architectures share a backbone and then branch into task-specific heads, while optimization techniques such as uncertainty weighting, GradNorm, or gradient surgery address conflicts between task gradients [126, 127, 128]. AiOIR, by contrast, treats each image degradation as a separate task but operates on the same input–output domain (i.e., pixel-level image reconstruction). Losses remain largely reconstruction-focused (perceptual loss, pixel loss), and task conflicts arise at the gradient level, often manifesting as mutual interference among degradation models. The design of the degradation-aware module is often insufficient to resolve conflicts among different degradations. The optimization process often receives less attention, leading to an oversight of the complex relationships and potential conflicts between multiple degradations in mixed training scenarios. Unlike unified models (e.g.,  [58, 105, 29, 54]) that employ mixed training, some novel studies address AiOIR from the perspective of MTL to resolve the inconsistencies and conflicts among different degradations, by treating each degradation as an independent task. By focusing on the optimization process and the interactions between tasks, these methods aim to mitigate conflicts and enhance overall performance. We can broadly classify AiOIR methods in MTL into two types: task grouping and task balancing.

### v3

@@ Abstract
Image restoration (IR) seeks to recover high-quality images from degraded observations caused by a wide range of factors, including noise, blur, compression, and adverse weather. While traditional IR methods have made notable progress by targeting individual degradation types, their specialization often comes at the cost of generalization, leaving them ill-equipped to handle the multifaceted distortions encountered in real-world applications. In response to this challenge, the all-in-one image restoration (AiOIR) paradigm has recently emerged, offering a unified framework that adeptly addresses multiple degradation types. These innovative models enhance the convenience and versatility by adaptively learning degradation-specific features while simultaneously leveraging shared knowledge across diverse corruptions. In this survey, we provide the first in-depth and systematic overview of AiOIR, delivering a structured taxonomy that categorizes existing methods by architectural designs, learning paradigms, and their core innovations. We systematically categorize current approaches and assess the challenges these models encounter, outlining research directions to propel this rapidly evolving field. To facilitate the evaluation of existing methods, we also consolidate widely-used datasets, evaluation protocols, and implementation practices, and compare and summarize the most advanced open-source models. As the first comprehensive review dedicated to AiOIR, this paper aims to map the conceptual landscape, synthesize prevailing techniques, and ignite further exploration toward more intelligent, unified, and adaptable visual restoration systems. A curated code repository is available at https://github.com/Harbinzzy/All-in-One-Image-Restoration-Survey.
@@ I Introduction
Image processing is an integral part of low-level vision tasks, and digital image processing has evolved significantly over the past few decades, transitioning from traditional methods to advanced deep learning techniques. These methods, while effective, are limited by their inability to handle complex and varied image degradation scenarios. With the rise of deep learning, image processing has achieved remarkable results, especially driven by convolutional neural networks (CNNs) [1] and Transformers [2]. In the image restoration (IR) field, single-task IR has achieved notable breakthroughs in correcting a specific type of image degradation (e.g., denoising [3, 4, 5, 6, 7, 8], dehazing [9, 10, 11, 12, 13], desnowing [14, 15, 16], deraining [17, 18, 19], deblurring [20, 21] and low-light image enhancement [22, 23, 24]). Although they have set state-of-the-art performance in restoring diverse degraded images, existing single-task IR methods generally lack the flexibility to adapt to new types of image degradation without extensive retraining (it is worth noting that some real-world IR methods [25, 26, 27, 28, 29] are not categorized as single-task IR). To alleviate the aforementioned issues, researchers have proposed all-in-one image restoration (AiOIR), devoting significant effort to introduce various key improvements (e.g., prompt-based learning approach, mixture-of-experts structure, multimodal model and autonomous agent) beyond single-task IR.
@@
With the rapid development of AiOIR, numerous researchers have collected a series of datasets tailored for various all-in-one task settings, e.g., the All-Weather dataset [30] for synthetic weather-related scenarios, the WeatherStream dataset [31] for real-world weather-related scenarios, the CDD-11 dataset [32] for composite degradations, etc. Leveraging these datasets, most recent works focus on improving the representation capability of IR networks for complicated degradation through well-designed methods based on prompt learning, contrastive learning, multimodal representations, etc. Although these works have achieved superior progress in the objective quality (e.g., PSNR, NIQE [33], and FID [34]), the restored images still suffer from unsatisfactory texture generation, hindering the practical deployment of IR methods in real-world scenarios. Collectively, these methodologies represent a substantial progression in the pursuit of sophisticated, accurate, and versatile AiOIR solutions.
@@
    Methodology (Sec. III): We provide a detailed analysis of representative AiOIR networks, aiming to illustrate the prevailing methods and elucidate the different categories of methods. By analyzing state-of-the-art methods, we summarize network architectures, learning strategies, some key improvements, and applications.
@@ II Preliminaries
In this section, we first define the concept of AiOIR and compare it with related concepts. Subsequently, we provide a comprehensive review of diverse scenarios in AiOIR and the corresponding approaches. Finally, we illustrate the interactions between different IR tasks.
@@ II-A Concept Definitions
Single-degradation image restoration (SDIR) focuses on recovering clean images from observations that have been corrupted by a specific type of degradation, such as noise [4, 5, 6], blur [20, 21], or rain [17, 18, 19]. These methods are typically task-specific, meaning that each model is designed for a particular degradation type. While these models demonstrate strong performance on known degradation types, they often exhibit limited generalization capability when encountering unseen degradation patterns or novel corruption intensities. Recently, some works [51, 52, 53] show certain generalizability to various degradation types with a unified framework. However, they need to train different models for different degradations, which are not all-in-one solutions as expected. We refer to these methods as multiple-degradation image restoration (MDIR). Subsequently, there have been recent attempts to unify multiple IR tasks within a single framework without the need for retraining. The seminal work [54] proposed the AiO adverse weather restoration method which utilizes a multi-encoder and decoder architecture and the neural architecture search across task-specific optimized encoders. Moreover, Li et al. [55] first proposed the all-in-one blind image restoration using contrastive learning to extract degradation representations, which can recover images from multiple corruptions without any degradation information in advance. The AiOIR approach aims to recover clean images under various conditions explicitly tailored to address isolated or mixed degradation based on multi-head and multi-tail structures, priors, or pretrained models within a unified framework. It is worth noting that some robust IR models [25, 27, 26, 28, 56, 57] are still broadly considered part of the AiOIR setting. The AiOIR models present practical advantages such as reduced storage requirements and simplified deployment. However, the principal challenge lies in developing robust architectures capable of effectively addressing diverse degradations and achieving high-quality restorations under varying conditions in real-world scenarios.
Figure 2: Milestones of AiOIR methods. We simply list their names and venues. The bar chart shows the approximate annual number of papers on AiOIR from 2020 to June 2025.
@@ II-B Typical Scenarios
We categorize AiOIR into several scenarios according to the types of images processed, including natural images, images under adverse weather conditions, underwater images, medical images, document images, hyperspectral images, as well as AI-generated images.
Natural Images. The degradation types commonly encountered in natural images include Gaussian noise, real-world noise, defocus blur, motion blur, low-light conditions, JPEG compression artifacts, mosaic effects, and issues arising from under-display cameras. Various AiOIR approaches have been developed to address these degradations, including techniques described in studies such as [58, 59, 60, 61, 62, 63]. These approaches primarily focus on resolving typical problems found in both natural and man-made scenes, enhancing the quality and usability of images affected by these common degradations.
Adverse Weather Conditions. In more extreme cases, the images requiring restoration may be severely impacted by various adverse weather conditions, such as snowflakes, raindrops, and dense haze. These conditions lead to ill-posed inverse problems that are crucial for applications in autonomous navigation, video retrieval, and outdoor surveillance. The AiOIR field has seen the emergence of many solutions aiming at addressing these challenges, including works like [64, 30, 65, 66, 67, 68]. Notably, Liu et al. [69] proposed multi-weather nighttime image restoration. These solutions are designed to effectively restore visibility and clarity in images affected by harsh environmental conditions.
Underwater Images. Underwater images suffer from wavelength-dependent light attenuation and scattering caused by marine microorganisms, leading to haze, color casts, low contrast, and loss of fine detail. UniUIR [57], a pioneering framework to tackle such compound distortions in an integrated manner, introduces the Mamba MoE module leveraging DepthAnythingV2 [70] to inject scene depth priors.
@@
Document Images. DocRes [73] pioneers the exploration of generalist models for AiOIR of document images. This work addresses five tasks (i.e., dewarping, deshadowing, appearance enhancement, deblurring, and binarization). DocRes employs a straightforward yet highly effective visual prompting method known as DTSPrompt, which effectively distinguishes between different tasks and accommodates various resolutions. TextDoctor [74], utilizing the structure pyramid prediction and patch pyramid diffusion model, exhibits superior performance in capturing the structure of inpainted text.
@@
AI-generated Images. With the rise of text-to-image (T2I) models, AI-generated images (AIGIs) often show inconsistent quality, creating an urgent need for models that better match human subjective ratings [77]. Li et al. proposed Q-Refine [78], a quality-aware refiner that enhances both the fidelity and the perceptual quality of AIGIs. G-Refine [79] further extends Q-Refine by simultaneously refining perceptual details and semantic alignment, achieving better generalization across various T2I models.
@@ II-C Task Interactions
These intertwined degradations present both challenges (task entanglement, ordering sensitivity, domain ambiguity) and opportunities (shared features, multi-task learning, unified representations) for AiOIR models. Understanding these interdependencies is essential for designing robust and generalizable AiOIR systems.
@@ III Methodology
All-in-one image restoration methods have gained significant attention for their ability to address multiple types of degradation within a unified model. This section analyzes AiOIR from four perspectives: (1) novel and traditional network architectures; (2) diverse learning strategies; (3) advanced techniques such as Prompt Learning, Mixture-of-Experts (MoE), Multimodal Models, and Agentic IR Systems and (4) applications of AiOIR. In Fig. 2, we present the representative works on AiOIR according to the timeline and method types.
@@ III-A Network Design for AiOIR
Existing AiOIR methods, as a form of multi-task learning (MTL), employ various architectural designs to handle inputs and outputs across multiple tasks, enabling efficient information sharing among them. Despite significant differences in their architectures, AiOIR methods can be broadly categorized into ten representative frameworks, as illustrated in Fig. 3 and described below.
@@
(f) Mixture-of-Experts Architecture: In MoE-based models, inputs are routed to different experts via a gating mechanism. For instance, MEASNet [86] considers both pixel-level and global features (including low-frequency and high-frequency components) to select the appropriate expert for IR.
@@
(h) Visual Prompting: These models employ a single encoder-decoder architecture and inject learnable visual prompts at multiple decoding stages to implicitly predict degradation conditions. Examples include works like [59, 93, 94, 95, 96]. The prompts guide the decoder to adaptively recovering various degraded images, serving as lightweight plug-and-play modules with minimal additional parameters.
@@ III-B Learning Strategies of AiOIR
Besides network design (as discussed in Sec. III-A), robust learning strategies are crucial for achieving satisfactory results in AiOIR. In this section, we discuss several promising learning strategies in this field. We begin with continual learning, exploring how it can prevent catastrophic forgetting. Next, we focus on contrastive learning and its application in complex degradation scenarios, emphasizing its role in enhancing the discriminative power of image features. We then discuss zero-shot learning in AiOIR, emphasizing its ability to generalize to unseen degradations. Additionally, we highlight the potential of multi-task learning (MTL) to optimize performance across various degradation tasks, pointing out the importance of addressing task relationships and conflicts. Finally, we introduce the concept of machine unlearning, exploring its potential for privacy protection.
@@ III-B2 Contrastive Learning
One of the significant challenges in IR is effectively handling unseen tasks and degradation types. The vast variability in possible degradations can severely impede a model’s generalization capabilities, making it less effective when confronted with new, unseen data. To address this issue, researchers have drawn inspiration from contrastive learning techniques that have proven to be successful in both high-level and low-level tasks [113, 114, 115]. Contrastive learning approaches typically serve as an additional form of regularization to improve the generalization of single-task restoration models [116, 117, 115, 118, 119]. By incorporating contrastive regularization, these methods aim to enhance model performance across a wide array of image restoration applications. In contrastive learning, the definitions of positive and negative samples can be flexibly adjusted, allowing researchers to tailor the learning process to better suit specific tasks and datasets. This flexibility ultimately enhances the model’s adaptability and performance in diverse image restoration scenarios [88]. Moreover, contrastive learning-based loss functions have been proposed to obtain discriminative degradation representations in the latent space [55, 58, 93, 120], further improving the AiOIR model’s ability to distinguish between different types of degradations and to generalize to unseen ones.
@@ III-B4 Multi-task Learning
MTL is a learning paradigm that leverages shared representations and knowledge across tasks, allowing models to learn more efficiently and effectively. By jointly learning from multiple objectives, MTL can improve generalization, reduce overfitting, and achieve better performance on individual tasks. Traditional MTL typically involves a combination of semantic tasks (e.g., classification, detection, segmentation). The tasks exhibit significant semantic differences and also possess distinct output spaces. Typical MTL architectures share a backbone and then branch into task-specific heads, while optimization techniques such as uncertainty weighting, GradNorm, or gradient surgery address conflicts between task gradients [130, 131, 132]. AiOIR, by contrast, treats each image degradation as a separate task but operates on the same input–output domain (i.e., pixel-level image reconstruction). Losses remain largely reconstruction-focused (perceptual loss, pixel loss), and task conflicts arise at the gradient level, often manifesting as mutual interference among degradation models. The design of the degradation-aware module is often insufficient to resolve conflicts among different degradations. The optimization process often receives less attention, leading to oversight of the complex relationships and potential conflicts between multiple degradations in mixed training scenarios. Unlike unified models (e.g.,  [59, 109, 30, 55]) that employ mixed training, some novel studies address AiOIR from the perspective of MTL to resolve inconsistencies and conflicts among different degradations, by treating each degradation as an independent task. By focusing on the optimization process and the interactions between tasks, these methods aim to mitigate conflicts and enhance overall performance. We can broadly classify AiOIR methods in MTL into two types: task grouping and task balancing.